# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

### v2

@@ Abstract
We present the second data release (DR2) of the Far-Infrared Polarimetric Large-Area CMZ Exploration (FIREPLACE) survey. This survey utilized the Stratospheric Observatory for Infrared Astronomy (SOFIA) High-resolution Airborne Wideband Camera plus (HAWC+) instrument at 214 µm (E-band) to observe dust polarization throughout the Central Molecular Zone (CMZ) of the Milky Way. DR2 consists of observations that were obtained in 2022 covering the region of the CMZ extending roughly from the Brick to the Sgr C molecular clouds (corresponding to a roughly 1 0.75 region of the sky). We combine DR2 with the first FIREPLACE data release covering the Sgr B2 region to obtain full coverage of the CMZ (a 1.5 0.75 region of the sky). After applying total and polarized intensity significance cuts on the full FIREPLACE data set we obtain 65,000 Nyquist-sampled polarization pseudovectors. The distribution of polarization pseudovectors confirms a bimodal distribution in the CMZ magnetic field orientations, recovering field components that are oriented predominantly parallel or perpendicular to the Galactic plane. These magnetic field orientations indicate possible connections between the previously observed parallel and perpendicular distributions. We also inspect the magnetic fields toward a set of prominent CMZ molecular clouds (the Brick, Three Little Pigs, 50 km s, Circum-nuclear Disk, CO 0.02-0.02, 20 km s, and Sgr C), revealing spatially varying magnetic fields that generally trace the morphologies of the clouds. We find evidence that compression from stellar winds and shear from tidal forces are prominent mechanisms influencing the structure of the magnetic fields observed within the clouds.
@@ 1 INTRODUCTION
The Central Molecular Zone (CMZ) of the Galactic Center (GC) is an extreme region of the Milky Way containing elevated molecular densities (10 – 10 cm; e.g., Mills et al., 2018) and magnetic field strengths (100s of G – mG, Yusef-Zadeh & Morris, 1987a; Chuss et al., 2003a; Pillai et al., 2015; Mangilli et al., 2019) compared to what is observed in the Galactic disk. A 3-color view of this region probed by radio and infrared observations is shown in Figure 1, where prominent molecular clouds throughout the CMZ have been labeled. The extreme properties of the CMZ, compared to those observed in the Galactic Disk, and the fact that it is the closest Galactic nuclear region to Earth, makes it an important region to study to deepen our understanding of galactic nuclear regions more generally.
@@
The star formation rate (SFR) is much lower in the CMZ than in the Galactic disk (e.g. Morris, 2023; Longmore et al., 2013). The low SFR is particularly unexpected given the amount of dense gas in the CMZ (e.g. 2-610 M, Barnes et al., 2017; Morris & Serabyn, 1996). There are multiple possible explanations for the low SFR in the CMZ, such as the strength and compressibility of turbulence in the region or that the CMZ could be in a period of inactivity between episodic bursts of star formation (Krumholz & Kruijssen, 2015). Another possibility could be the elevated magnetic field strengths observed in the region.
@@
The role of the magnetic field in the CMZ has been studied using both radio and infrared observations. Radio observations reveal a population of synchrotron structures known as non-thermal filaments (NTFs) that are understood to trace ordered magnetic fields. These NTFs are predominantly oriented perpendicular to the Galactic plane (Yusef-Zadeh et al., 2022; Heywood et al., 2022; Morris, 1996). The general orientation of the NTF population has been interpreted to indicate a large-scale poloidal magnetic field that pervades the GC (Morris, 2006; Yusef-Zadeh et al., 2004).
@@
The magnetic field derived from studies of dust polarization at sub-millimeter and infrared wavelengths, conversely, reveal a magnetic field that seemingly conflicts with the poloidal magnetic field traced by the NTFs. The magnetic field derived from the dust polarization observations traces the most extended axes of the molecular clouds within the CMZ – an orientation that is largely parallel to the Galactic plane and consistent with a toroidal field (Guan et al., 2021; Mangilli et al., 2019; Nishiyama et al., 2010; Novak et al., 2000). It is unclear how the seemingly distinct magnetic fields traced by the infrared and radio observations are related; however, observations at 90 and 150 GHz reveal regions where the magnetic field becomes poloidal rather than toroidal (Guan et al., 2021). This could either indicate connections between the fields traced by dust polarization and the NTFs or a spatially variable distribution comprised of multiple magnetic field components (Guan et al., 2021). The latter is likely given the sensitivity of these observations to both thermal dust and synchrotron emission.
@@
The Far-Infrared Polarimetric Large Area CMZ Exploration (FIREPLACE) survey was proposed to address the question of how the magnetic field at infrared wavelengths varies throughout the CMZ by utilizing observations made by the High-resolution Airborne Wideband Camera plus (HAWC+) instrument on the Stratospheric Observatory for Infrared Astronomy (SOFIA). The first FIREPLACE data release (DR1) covered a region of the CMZ centered on Sgr B2 (Butterfield et al., 2023, hereafter reffered to as FIREPLACE I). The field of view of these first FIREPLACE observations is marked by the green dashed rectangle in Figure 1. The DR1 observations reveal a more spatially varying magnetic field at 19.6″ resolution than is seen in previous CMZ-wide surveys (1′, Guan et al., 2021; Mangilli et al., 2019). An in-depth study of the magnetic field system local to one of the molecular structures identified in DR1 indicates that the magnetic field probed by FIREPLACE traces smaller-scale molecular structures local to the CMZ than in previous CMZ-wide surveys (Butterfield et al., 2024, hereafter referred to as FIREPLACE II).
@@
In this paper we present the second FIREPLACE data release (DR2), which is comprised of SOFIA/HAWC+ observations covering the CMZ from the Brick to Sgr C (a roughly 1 0.75 region of the sky). We then combine the DR1 and DR2 FIREPLACE data sets to obtain a polarimetric map with full coverage of the CMZ (a 1.5 0.75 region). In Section 2 we describe the data-analysis procedure applied to the DR2 FIREPLACE observations. In Section 3 we describe the data quality checks performed to verify the reduction procedure used to make maps from the raw observations obtained from SOFIA/HAWC+. The CMZ-wide FIREPLACE observations are then presented in Section 4, where the reduced DR1 and DR2 data sets are combined. In Section 5 we study the detailed polarimetric properties of prominent molecular clouds in our DR2 data set and we present our conclusions in Section 6.
@@ 2.1 DR2 Observations: 2022 SOFIA Flights
FIREPLACE DR2 is an expansion of the FIREPLACE pilot program. The DR1 pilot program observed the eastern third of the CMZ (the greater Sgr B2 complex) during SOFIA flights F775 and F777 on August 31, 2021 and September 2 2021, respectively. This DR2 release incorporates observations obtained from several 2022 SOFIA flights: four Southern hemisphere Christchurch-based flights in June and July 2022 (flights F890, F891, F893, and F895) and three Northern hemisphere Palmdale-based flights in September 2022 (flights F916, F917, and F918).
@@
The full FIREPLACE survey covers the central 1.5 0.75 of the CMZ. This is the largest polarimetric map made by SOFIA. Given both the large map size and the extended flux in this region at 214 µm, the standard Chop-Nod-Match (CNM Harper et al., 2018) observing strategy was not practical. Instead, FIREPLACE utilized the on-the-fly mapping (OTFMAP) mode. The OTFMAP mode involves rapid scanning of the array over the field to create time streams that are subsequently fit for correlated noise components and map values using the SOFIA data reduction pipeline, which is based on the CRUSH algorithm (Kovács, 2008). This is done for each of four angular positions of the half-wave plate (0, 22.5, 45, and 67.5), and the Stokes , , and parameters are fit to the resulting images.
@@
Even with the OTFMAP mode strategy, the full extent of the CMZ is larger than what can be feasible observed with a single scan. Therefore we developed an observing plan utilizing a set of fields that fully cover the CMZ. The fields are reduced using the SOFIA OTFMAP reduction software as described above and then merged into a final map. This observation approach is novel, and there is significant correlated noise on short timescales. Therefore, care has to be taken in optimizing the observations and verifying the data set.
@@
For FIREPLACE, each scan is designed to cross the Galactic plane at an angle of either 30 relative to Galactic North. For each scan, we follow the same parameters as FIREPLACE I – the boresight of the detector follows a Lissajous scan with amplitudes of 27′ 4.5′. The ratio of the frequencies between the parametric sinusoidal curves of the Lissajous pattern is set to , with the higher frequency corresponding to the long axis. The total time for each scan is 120 seconds, and the scan rate is 300″ s. The planned boresight path for each field of view is shown in Figure 2. The numbering in the figure marks the center of each field and the area covered by each boresight trajectory is indicated by the semi-transparent rectangles. Each field was observed using two different orientations that are oriented 30 with respect to Galactic North. The region targeted for polarimetry measurements is shown by the white contour (at a level of 7189 MJy sr at 250 µm). Each scan was designed with two considerations in mind. First, we required the scan to have sufficient extent so that on each scan the detector array crosses an effective “zero reference,” where the intensity is significantly lower than the target region. Second, we required significant overlap between adjacent and crossing fields for repeated measurements of any sky position in our target region. This latter point enables sufficient cross-linking between map pixels and detectors for the simultaneous fitting of all of the time streams for the map. Although the current version of the SOFIA pipeline does not allow for such simultaneous fitting, future data reduction efforts are being explored to take advantage of this aspect of the observation plan for increased fidelity, especially on larger spatial scales.
@@
Observations in the region spanning fields 1 - 5 were taken in 2021 and comprise the DR1 FIREPLACE observations presented in FIREPLACE I. Note that the 2021 fields were implemented differently than those shown in Figure 2 due to setup issues during the 2021 observation runs (see Figure 2 of FIREPLACE I which shows the orientations of the DR1 observation fields). DR2 re-observed some of the fields covered in DR1 (fields 4 and 5), to increase the spatial overlap between the DR1 and DR2 observations, but primarily focuses on fields 6-18 to complete the map of the CMZ (Figure 2).
@@ 2.2 Data Reduction
To reduce the raw observations obtained from the SOFIA Infrared Science Archive at IPAC (IRSA) we use the HAWC+ Data Reduction Pipeline (DRP). We used DRP version 2.7.0, which is a Python-wrapper built on a Java version of the CRUSH reduction algorithm (Kovács, 2008). Since the latest version of of DRP (version 3.2.0) has not been verified for such large maps with extended emission we did not use this more recent version of DRP to reduce our HAWC+ observations. The use of DRP version 2.7.0 has the added benefit of maintaining consistency with DR1, which also used this DRP version.
@@
To reduce these observations using CRUSH we set the rounds parameter to “-rounds=85,” which increases the number of iterations CRUSH performs when numerically modeling the correlated noise. We also set the “-extended” option, which indicates that CRUSH needs to account for extended emission during data reduction to better preserve extended source emission when identifying correlated noise. Furthermore, we specified “-fixjumps” and “-downsample=1” to remove the effects of quantum flux jumps in the SQUID amplifiers and to prevent spatial averaging, respectively. These CRUSH settings match the parameters used for the FIREPLACE DR1 data release presented in FIREPLACE I.
@@
Using these parameters in CRUSH, we reduced each field separately. For each field all scans were analyzed simultaneously where possible (8 scans for most fields accounting for 2 sets of 4 HWP angles). In some cases, an observation consisted of only one set of 4 scans. In these cases, 4 scans were included in the CRUSH step. Table A.2.1 in Appendix A.1 provides details on each of the scans observed and their mapping to the fields shown in Figure 2. Table A.2.1 also indicates how many scans were performed for each field.
@@
Due to systematically high correlated noise that was evident in the post-reduction scans for the relevant fields, data from flight F890 were omitted from the final reduction. Fortunately, the regions observed during this flight were also observed during other flights, and so we still obtain full CMZ coverage. For completeness Table A.2.1 includes information on the observations conducted during flight F890, but we emphasize that the results presented in this work do not incorporate the observations obtained from that flight. The uncalibrated observations can be obtained from the SOFIA:IRSA data archive111https://irsa.ipac.caltech.edu/frontpage/ and can be found using the proposal ID of the survey (09_0054).
@@
After reducing each field using CRUSH, the resulting total intensity emission of the fields was cross-correlated with Herschel 250 µm observations from Molinari et al. (2011) to identify pointing offsets in our fields. We found that the majority of fields corresponding to the 2022 observations had pointing offsets on the scales of 10″(corresponding to offsets 2 or more pixels in our map). These significant pointing offsets were corrected in all fields in which an offset of more than 5″ (angular size of 1 pixel in our map) was identified. Correcting the FIREPLACE pointings using the Herschel 250 µm observations completely corrects these offsets. Offsets smaller than 5″ are below the pixel size of our observations, so we did not implement offset corrections for these fields.
@@
After applying the offset correction, the instrumental polarization is then removed from each field. The calibrated, offset-corrected, and instrumental polarization-corrected fields were then co-added into one combined data set using the DRP merge algorithm. Figure 3 shows the total intensity distribution for the merged 2022 observations with polarization pseudovectors shown as yellow lines. These polarization angles are calculated using
@@
where and are defined following the IAU convention where North corresponds to an angle of 0 with increasing in the counterclockwise direction222The original reduction was done in equatorial coordinates. Thus “North” here refers to “equatorial North.” When switching to Galactic coordinates, the angle is rotated to be relative to “Galactic North.”.
@@
where the estimate of the uncertaintainty of the measured fractional polarization propagated from the fitted weights in the CRUSH step, and is the measured fractional polarization. Pseudovectors are only derived for lines of sight having total intensity above a significance threshold of , percentage polarizations less than 50% ( 50%), and significant polarized intensity (), which are quality cuts consistent with standard SOFIA polarimetry practice (Gordon et al., 2018).

### v3

@@ Abstract
We present the second data release (DR2) of the Far-InfraREd Polarimetric Large-Area CMZ Exploration (FIREPLACE) survey. This survey utilized the Stratospheric Observatory for Infrared Astronomy (SOFIA) High-resolution Airborne Wideband Camera plus (HAWC+) instrument at 214 µm (E-band) at a resolution of 19.6″ to observe thermal polarized dust emission throughout the Central Molecular Zone (CMZ). DR2 consists of observations obtained in 2022 covering the region of the CMZ extending from the Brick to the Sgr C molecular clouds (corresponding to a 1∘ 0.75∘ region of the sky). We combine DR2 with the first FIREPLACE data release (DR1) to obtain full coverage of the CMZ (a 1.5∘ 0.75∘ region of the sky). After applying total and polarized intensity significance cuts on the full FIREPLACE data set, we obtain 64,000 Nyquist-sampled polarization pseudovectors. The distribution of polarization pseudovectors confirms a bimodal distribution in the CMZ magnetic field orientations, recovering field components that are oriented predominantly parallel or perpendicular to the Galactic plane. This distribution of orientations is similar to what was observed in DR1 and other studies. We also inspect the magnetic fields toward a set of prominent CMZ molecular clouds (the Brick, Three Little Pigs, 50 km s-1 and 20 km s-1 clouds, Circum-nuclear Disk, CO 0.02-0.02, and Sgr C), revealing spatially varying magnetic fields having orientations that generally trace the total intensity morphologies of the clouds. We find evidence that compression from stellar winds and shear from tidal forces are prominent mechanisms influencing the structure of the magnetic fields.
@@ 1 INTRODUCTION
The Central Molecular Zone (CMZ) of the Galactic Center (GC) is an extreme region of the Milky Way, containing elevated molecular densities (103 – 106 cm-3; e.g., Mills et al., 2018) and magnetic field strengths (100s of G – mG, Yusef-Zadeh & Morris, 1987a; Plante et al., 1995; Chuss et al., 2003a; Pillai et al., 2015; Hsieh et al., 2018; Mangilli et al., 2019; Guerra et al., 2023) compared to what is observed in the Galactic disk. A 3-color view of this region probed by radio and infrared observations is shown in Figure 1, where prominent molecular clouds throughout the CMZ have been labeled. The combination of its extreme properties and proximity to Earth compared to other galactic nuclear regions makes it a useful proxy for more distant gas-rich galactic nuclei throughout the universe.
@@
The star formation rate (SFR) is much lower in the CMZ than in the Galactic disk (e.g. Longmore et al., 2013; Morris, 2023). The low SFR is particularly unexpected given the amount of dense gas in the CMZ (e.g. 2-6107 M⊙, Morris & Serabyn, 1996; Barnes et al., 2017). There are multiple possible explanations for the low SFR in the CMZ, such as the strength and compressibility of turbulence in the region or that the CMZ could be in a period of inactivity between episodic bursts of star formation (Krumholz & Kruijssen, 2015). Another factor that could contribute to the inhibition of star formation in the CMZ is the elevated magnetic field strengths observed in the region (e.g., Morris, 1989, 1993).
@@
The distribution of the magnetic field in the CMZ has been studied using both radio and infrared observations. Radio observations reveal a population of synchrotron structures known as non-thermal filaments (NTFs) that are understood to trace ordered magnetic fields. These NTFs are predominantly oriented perpendicular to the Galactic plane (Morris, 1996; Yusef-Zadeh et al., 2022; Heywood et al., 2022). The general orientation of the NTF population has been interpreted to indicate a large-scale vertical (or possibly poloidal) magnetic field that pervades the GC (Yusef-Zadeh et al., 2004; Morris, 2006).
@@
Conversely, the magnetic field derived from large-scale studies of dust polarization at sub-millimeter and infrared wavelengths reveals a distribution that seemingly conflicts with the vertical magnetic field traced by the NTFs. The magnetic field derived from the dust polarization observations is aligned parallel to the distribution of the molecular clouds within the CMZ – an orientation that is largely parallel to the Galactic plane and consistent with a horizontal (or possibly toroidal) field (Novak et al., 2000; Nishiyama et al., 2010; Mangilli et al., 2019; Guan et al., 2021).
One possible explanation for these different magnetic field orientations could be the distinct mechanisms responsible for the polarization (relativistic electrons illuminating the NTFs for the radio observations compared to dust-grain alignment for the infrared observations). However, polarimetric observations at 90 and 150 GHz reveal regions where the magnetic field becomes vertical rather than horizontal (Guan et al., 2021). These vertical magnetic field regions could either indicate connections between the fields traced by dust polarization and the NTFs or a spatially variable distribution comprised of multiple magnetic field components (Guan et al., 2021). The latter is likely given the sensitivity of these particular observations to both thermal dust and synchrotron emission.
@@
The Far-InfraREd Polarimetric Large Area CMZ Exploration (FIREPLACE) survey was proposed to address the question of how the magnetic field at infrared wavelengths varies throughout the CMZ by utilizing observations made by the High-resolution Airborne Wideband Camera plus (HAWC+) instrument on the Stratospheric Observatory for Infrared Astronomy (SOFIA). The first FIREPLACE data release (DR1) covered a region of the CMZ centered on Sgr B2 (Butterfield et al., 2024a, hereafter reffered to as FIREPLACE I). The field of view of these first FIREPLACE observations is marked by the green dashed rectangle in Figure 1. The DR1 observations reveal a more spatially varying magnetic field at 19.6″ resolution than is seen in previous CMZ-wide surveys (1′, Mangilli et al., 2019; Guan et al., 2021). An in-depth study of the magnetic field system local to one of the molecular structures identified in DR1 indicates that the magnetic field probed by FIREPLACE traces smaller-scale molecular structures local to the CMZ than in these previous CMZ-wide surveys (Butterfield et al., 2024b, hereafter referred to as FIREPLACE II).
@@
In this paper we present the second FIREPLACE data release (DR2), which is comprised of SOFIA/HAWC+ observations covering the extent of the CMZ from the Brick to Sgr C (a roughly 1∘ 0.75∘ region of the sky). We then combine the DR1 and DR2 FIREPLACE data sets to obtain a polarimetric map with full coverage of the CMZ (a 1.5∘ 0.75∘ region). In Section 2 we describe the data reduction procedure applied to the DR2 FIREPLACE observations to make maps from the raw observations obtained from SOFIA/HAWC+. Section 3 describes the data quality checks performed to verify the reduction procedure. The CMZ-wide FIREPLACE observations are then presented and discussed in Section 4. In Section 5 we study the detailed polarimetric properties of prominent molecular clouds in our DR2 data set and we present our conclusions in Section 6.
@@ 2.1 DR2 Observations: 2022 SOFIA Flights
FIREPLACE DR2 is an expansion of the FIREPLACE pilot program (FIREPLACE I). The DR1 pilot program observed the eastern third of the CMZ (the greater Sgr B2 complex) during SOFIA flights F775 and F777 on August 31, 2021 and September 2 2021, respectively. The new DR2 release expands on this pilot program to include observations obtained from several 2022 SOFIA flights: four Southern hemisphere Christchurch-based flights in June and July 2022 (flights F890, F891, F893, and F895) and three Northern hemisphere Palmdale-based flights in September 2022 (flights F916, F917, and F918).
@@
The full FIREPLACE survey covers the central 1.5∘ 0.75∘ of the CMZ with a pixel size of 5.3″, beam size of 19.6″, central wavelength of 214 µm, and a bandwidth of 175 µm. This is the largest polarimetric map made by SOFIA. Given both the large map size and the extended flux in this region at 214 µm, the standard Chop-Nod-Match (CNM, Harper et al., 2018) observing strategy was not practical. Instead, FIREPLACE utilized the on-the-fly mapping (OTFMAP) mode. The OTFMAP mode involves rapid scanning of the HAWC+ array over the field to create time streams that are subsequently fit for correlated noise components and map values using the SOFIA data reduction pipeline, which is based on the CRUSH algorithm (Kovács, 2008). This is done for each of four angular positions of the half-wave plate (0∘, 22.5∘, 45∘, and 67.5∘), and the Stokes , , and parameters are fit to the resulting images. A detailed description of the OTFMAP polarimetric observing mode using HAWC+ is presented in Lopez-Rodriguez et al. (2022).
@@
Even with the OTFMAP mode strategy, the full extent of the CMZ is larger than what can be feasibly observed with a single scan. Therefore we developed an observing plan utilizing a set of fields that fully cover the CMZ as shown in Figure 2. The fields are reduced individually using the SOFIA OTFMAP reduction software as described below and then merged into a final map. This observation approach is novel, and there is significant correlated noise on short timescales. Therefore, care has to be taken planning the observations, in optimizing the reduction parameters, and verifying the data set. The reduction process is described in Section 2.2 and the data verification methods are discussed in Section 3.
@@
For FIREPLACE, each scan is designed to cross the Galactic plane at an angle of either 30∘ relative to Galactic North. For each scan, we follow the same parameters as FIREPLACE I – the boresight of the detector follows a Lissajous scan with amplitudes of 27′ 4.5′. The ratio of the frequencies between the parametric sinusoidal curves of the Lissajous pattern is set to , with the higher frequency corresponding to the long axis. The on-source time for each scan is 120 seconds and the scan rate is 300″ s-1. The planned boresight path for each field of view is shown in Figure 2. The numbering in the figure marks the center of each field and the area covered by each boresight trajectory is indicated by the semi-transparent rectangles. Each field was observed using two different scan alignments that are oriented 30∘ with respect to Galactic North. The region targeted for polarimetry measurements is shown by the white contour (at a level of 7189 MJy sr-1 at 250 µm).
OTFMAP mode is inherently a differential image reconstruction technique that does not preserve the map intensity zero point. For this reason, each scan was designed with two considerations in mind. First, scans were required to have sufficient extent for the detector array to cross an effective “zero reference” common to other scans, where the instrument field of view encounters a low-intensity reference region that is subdominant to CMZ molecular cloud intensity. Second, an overlap of two-thirds is employed between adjacent fields for repeated measurements of any sky position in our target region. We used the same overlap for both scan orientations to ensure additional measurements and improve the cross-linking of our observations. This latter point enables sufficient cross-linking between consecutive scans for the simultaneous fitting of all of the time streams when generating the combined map.
Although the current version of the SOFIA pipeline does not allow for simultaneous fitting over all observational scans, future data reduction efforts are being explored to take advantage of this aspect of the scan strategy for increased fidelity in recovering emission from larger-scale structures. This future analysis can incorporate CMB map-making strategies to reconstruct spatial structures larger than the instantaneous array field of view. Full sky implementations have been explored previously (e.g., Wright et al., 1996; Tegmark, 1997). More recently there are examples of implementations at smaller angular scales (Aiola et al., 2020; Qu et al., 2023; Eimer et al., 2023; Li et al., 2023).
@@
Observations in the region spanning fields 1 - 5 were taken in 2021 and comprise the DR1 FIREPLACE observations presented in FIREPLACE I. Note that the 2021 fields were implemented differently than those shown in Figure 2 due to setup issues during the 2021 observation runs (see Figure 2 of FIREPLACE I which shows the actual orientations of the DR1 observation fields). DR2 re-observed some of the fields covered in DR1 (fields 4 and 5), to increase the spatial overlap between the DR1 and DR2 observations. However, DR2 primarily focuses on fields 6-18 to complete the map of the CMZ (Figure 2).
@@ 2.2 Data Reduction
To reduce the raw observations obtained from the SOFIA Infrared Science Archive at IPAC (IRSA)111https://irsa.ipac.caltech.edu/frontpage/ we use the HAWC+ Data Reduction Pipeline (DRP). We used DRP version 2.7.0, which is a Python-wrapper built on a Java version of the CRUSH reduction algorithm (Kovács, 2008). Since the latest version of DRP (version 3.2.0) has not been verified for such large maps with extended emission we did not use this more recent version of DRP to reduce our HAWC+ observations. The use of DRP version 2.7.0 has the added benefit of maintaining consistency with DR1, which also used this DRP version.
@@
To reduce the DR2 observations using CRUSH we set the rounds parameter to “-rounds=85,” which increases the number of iterations CRUSH performs when numerically modeling the correlated noise. We also set the “-extended” option, which indicates that CRUSH needs to account for extended emission during data reduction to better preserve large scale flux when identifying correlated noise. Furthermore, we specified “-fixjumps” and “-downsample=1” to remove the effects of quantum flux jumps in the SQUID amplifiers and to prevent spatial averaging, respectively. These CRUSH settings match the parameters used for the FIREPLACE DR1 data release presented in FIREPLACE I. For each reduced field, DRP removes the 2% instrumental polarization that Harper et al. (2018) found in all HAWC+ bands.
Previous polarimetric studies using SOFIA/HAWC+ observations have found that using the “-extended” flag introduced artificial polarization in low-intensity regions (Lopez-Rodriguez et al., 2022; Le Gouellec et al., 2023). We observed this same phenomenon when reducing the FIREPLACE observations using DRP version 3.2.0, but not when using DRP version 2.7.0. Figure A.1 in Appendix A.1 provides a representative example of the artifacts encountered when using DRP version 3.2.0 for a single FIREPLACE field. The data reductions presented in Le Gouellec et al. (2023) and Lopez-Rodriguez et al. (2022) also describe an extensive “zero-level background” correction. In this work we use the filtered version of the FIREPLACE observations obtained from the CRUSH output and do not correct for the intensity “zero-point.”
@@
Using the parameters specified above we reduced each field separately. For each field, all scans were analyzed simultaneously where possible (8 scans for most fields accounting for 2 sets of 4 HWP angles). In some cases, an observation consisted of only one set of 4 scans. In these cases, 4 scans were included in the CRUSH step. Table A.1 in Appendix A.2 provides details on each of the scans observed and their mapping to the fields shown in Figure 2. Table A.1 also indicates how many scans were obtained for each field.
@@
Due to systematically high correlated noise that was evident in the post-reduction scans for the relevant fields, data from flight F890 were omitted from the final reduction. Fortunately, the regions observed during this flight were also observed during other flights, and so we still obtain full CMZ coverage. For completeness, Table A.1 includes information on the observations conducted during flight F890, but we emphasize that the results presented in this work do not incorporate the observations obtained from that flight.
@@
After reducing each field using CRUSH, the resulting total intensity emission of the fields was cross-correlated with Herschel 250 µm observations from Molinari et al. (2011) to identify pointing offsets in our fields. We found that the majority of fields corresponding to the 2022 observations had pointing offsets on the scales of 10″(corresponding to offsets of 2 or more pixels in our map). These significant pointing corrections were applied to all fields in which an offset of more than 5″ (angular size of 1 pixel in our map) was identified. Correcting the FIREPLACE pointings using the Herschel 250 µm observations completely corrects these offsets. Offsets smaller than 5″ are below the pixel size of our observations, so we did not implement offset corrections for these fields.
After applying the offset correction, the instrumental polarization is then removed from each field. The DRP has a built-in instrumental polarization correction which has been applied to the reduced FIREPLACE observations. The instrumental polarization is determined using images of the sky, and the polarization from these calibrated observations is subtracted from the and FIREPLACE data sets. This correction is expected to be accurate within / and / 0.3% according to the HAWC+ handbook222https://irsa.ipac.caltech.edu/data/SOFIA/docs/instruments/hawc/index.html.
@@
The calibrated, offset-corrected, and instrumental polarization-corrected fields were then co-added into one combined data set using the DRP merge algorithm. The merge algorithm resamples the input fields onto a common pixel grid based on the input World Coordinate Systems of the individual fields. The input pixels are then weighted based on their distance from the nearest coordinate in the common grid. Figure 3 shows the total intensity distribution for the merged 2022 observations with polarization angle pseudovectors shown as yellow lines. These polarization angles are calculated using
@@
where and are defined following the IAU convention where North corresponds to an angle of 0∘ with increasing in the counterclockwise direction333The original reduction was done in equatorial coordinates. Thus “North” here refers to “equatorial North.” When switching to Galactic coordinates, the angle is rotated to be relative to “Galactic North.”.
@@
where is the estimate of the uncertaintainty of the measured fractional polarization propagated from the fitted weights in the CRUSH step, and is the measured fractional polarization. Pseudovectors are only derived for lines-of-sight having total intensity above a significance threshold of , percentage polarizations less than 50% ( 50%), and significant polarized intensity (), which are quality cuts consistent with standard SOFIA polarimetry practice (Gordon et al., 2018). The individual calibrated files and the merged map of the entire CMZ are publicly available and can be downloaded from the NASA Infrared Processing and Analysis Center (IPAC) Legacy Surveys webpage as FITS files444https://irsa.ipac.caltech.edu/data/SOFIA/docs/data/legacy-programs/two-color-polarimetric-survey-galactic-center-pilot-legacy-program/index.html.